The length of a rectangle is 3m less than twice the width and the area of the rectangle is 44m. find the dimensions of the rectangle
Answer by josgarithmetic(39617)

L=-3+2w, and wL=44.
Substitute: w(2w-3)=44.
Simplify and solve.